OpenStudy (anonymous):

A 47-foot-long piece of pipe is to be cut into three pieces. The second piece is three times as long as the first piece, and the third piece is two feet more than five times the length of the first piece. What is the length of the longest piece? 5 feet 15 feet 27 feet 32 feet

Mehek (mehek14):

47 = 3x + (2+5x) solve for x

Mehek (mehek14):

nw substitute any value of the answer choices above and see which one is correct

Mehek (mehek14):

lets start of by substituting 5 for x 47 = 3(5) + (2+5(5)) 47 = 15 + (2 + 25) 47 = 15 + 27 47 = 47 so the first answer choice is correct